Uniqueness of the coexistence state of predator-prey systems on the whole ℝ

Rosa Pardo Departamento de Matematica Aplicada Facultad de Ciencias Quimicas Universidad Complutense de Madrid 28040 Madrid Spain

                rpardo@sunma4.mat.ucm.es

It is shown that predator-prey systems on the whole real line have a unique coexistence state. To obtain this result, we proof the invertibility of certain noncooperative elliptic system.